### v1

# Dynamic Level Sets

Abstract
A mathematical concept is identified and analyzed that is implicit in the 2012 paper Turing Incomputable Computation [6], presented at the Alan Turing Centenary Conference (Turing 100, Manchester). The concept, called dynamic level sets, is distinct from mathematical concepts in the standard literature on dynamical systems, topology, and computability theory. A new mathematical object is explained and why it may have escaped prior characterizations, including the classical result of de Leeuw, Moore, Shannon, and Shapiro [4] that probabilistic Turing machines compute no more than deterministic ones.
1 Level Sets in Pure Mathematics
In standard analysis, topology, and dynamical systems theory, a level set of a real-valued function $f:X\to\mathbb{R}$ at value $c$ is the preimage
$f^{-1}(c)\;=\;\{\,x\in X\mid f(x)=c\,\}.$
Closely related are the sublevel sets $f^{-1}((-\infty,c])=\{x\mid f(x)\leq c\}$.
While contour lines [^1] Charles Hutton used them in his 1774 survey of Schiehallion mountain in Scotland. on maps [10] (curves of equal elevation) represent the earliest graphical use of the concept, the first systematic mathematical use of level sets in dynamics appears in Lyapunov’s 1892 doctoral dissertation [12]. Lyapunov’s stability analysis depends on the geometry of sublevel sets of a positive-definite function $V$: the invariance of these sets under the flow verifies stability, while their outward drift proves instability.
Level sets in classical dynamical systems are static geometric objects. The phase space itself, where level sets reside, was co-developed in the late nineteenth century by Boltzmann [3], Poincaré [13], and Gibbs [11]. Boltzmann’s 1872 paper introduced the distribution function on position-momentum space; Poincaré’s four-part memoir of 1881–1886 established the qualitative geometric theory of trajectories in the plane; Gibbs’s 1902 treatise systematized the ensemble formalism. In this framework, the level sets of any function associated with a dynamical system $\dot{x}=F(x)$ (conserved quantities, Lyapunov functions, or the vector field itself) are fixed structures. Trajectories move through them, but the level sets do not move.
The entire geometric theory of dynamical systems (invariant manifolds, stable and unstable manifolds, basins of attraction, separatrices) rests on this assumption, as codified in the foundational survey of Smale [15] and in the Russian school’s treatment of bifurcations [1, 2]. In the latter, bifurcations alter the topology of level sets at isolated parameter values, but always according to equations fixed in advance; between bifurcation points the level set structure is entirely permanent.
Remark 1.1.
This static property is not a technical limitation, but a foundational assumption. A classical dynamical system is a triple $(T,X,\Phi)$ where $\Phi:X\times T\to X$ is a fixed evolution rule. The level sets of any function associated with the system are therefore determined once and for all when $\Phi$ is specified.
2 Level Sets in the Active Element Machine
The Active Element Machine (AEM) [5], when executing a Universal Turing Machine (UTM) program $\eta$, physically realizes the boolean transition functions
$\eta_{k}:\{0,1\}^{3}\times\{0,1\}^{2}\;\longrightarrow\;\{0,1\},\quad k\in\{0,1,2,3,4,5\},$
as active element firing patterns. Each function $\eta_{k}$ has a fixed logical level set structure. For example, $\eta_{3}^{-1}\{1\}\;=\;$
$\bigl\{\,(111,00),\,(110,00),\,(110,01),\,(110,10),\,(101,00),\,(101,01),\,(101,10),$
$(100,00),\,(100,10),\,(011,01),\,(011,10),\,(010,11),\,(010,01),\,(010,00)\bigr\}$,
which is a property of the Minsky UTM program being executed.
What is not fixed is the physical realization of these level sets within the AEM. At each UTM computational step, quantum random bits $R_{0},R_{1},\ldots,R_{13}$ are generated by a quantum random number generator of the type described in [16]; see also [9] for a comprehensive review of such devices. Meta commands set_dynamic_C and set_dynamic_E then reconfigure the active elements $D_{0},\ldots,D_{13}$ and their connections (the physical objects that compute membership in $\eta_{3}^{-1}\{1\}$), based on which of the $R_{k}$ fired. The same logical level set is realized by a different spatio-temporal firing pattern at each computational step.
More precisely, for each UTM instruction $(q,\alpha)\in Q\times A$, an invertible boolean function is defined
$B_{(q,\alpha)}:\{0,1\}^{15}\;\longrightarrow\;\{0,1\}^{15}$
(shown in Table 4 of [6]) that maps the random firing pattern $(x_{0},\ldots,x_{13},b_{3})$ to the firing activity $(y_{0},\ldots,y_{13},p_{3})$ of the elements computing $\eta_{3}$. The choice of which $B_{(q,\alpha)}$ governs the realization at step $j$ depends on the current UTM instruction $I_{j}$ – that is, on the dynamical state of the computation itself.
3 A New Mathematical Object
The concept is stated more precisely.
Definition 3.1 (Dynamic Level Set).
Let $f:X\to\{0,1\}$ be a fixed boolean function with level sets $f^{-1}\{0\}$ and $f^{-1}\{1\}$. A dynamic level set decomposition of $f$ is a family of invertible boolean functions $\{B_{s}\}_{s\in S}$, indexed by a state space $S$, together with an incomputable sequence $\omega:\mathbb{N}\to\{0,1\}^{n}$ and a computable function $h:\mathbb{N}\to S$, such that at step $j$ the physical encoding of $f^{-1}\{1\}$ is determined by $B_{h(j)}(\omega(j))$ rather than by any fixed representation. The level set as a mathematical set is invariant; its physical instantiation is a dynamical variable.
Remark 3.1.
This captures what is constructed in Procedure 2  [6]. The logical level sets $\eta_{k}^{-1}\{1\}$ are invariant properties of the UTM program $\eta$. The physical realizations are dynamical variables of the AEM, evolving by a rule that depends on the current UTM instruction (computable via $h$) and on quantum random input (incomputable by the two quantum randomness axioms; see [16, 9]).
A dynamic level set is novel for the following reasons.
It is not a parametric family.
In a parametric family $F(x;\,\theta(t))$, the level sets $F^{-1}(c;\,\theta(t))$ sweep through a predetermined sequence as $\theta$ varies. Here, no predetermined sequence exists: the quantum randomness axioms guarantee that the sequence of level set realizations is itself Turing incomputable, so it cannot have been fixed in advance.
It is not bifurcation theory.
In bifurcation theory [1, 2], level set topology changes at isolated parameter values according to fixed equations of motion. In procedure 2  [6] , the physical realization changes at every computational step, not at exceptional points, and is governed by an incomputable process rather than a smooth family.
It is not a probabilistic Turing machine in the sense of de Leeuw et al.
The classical 1956 result shows that a Turing machine augmented with a random oracle with probability $p$, where $p$ is Turing computable, computes no more than a deterministic one. That result applies because the Turing machine program, and therefore its level set structure, does not change in response to random input. In the AEM, the Meta command reconfigures the physical realization of the level sets at each step based on the random input. The program’s logical content is invariant; its physical encoding is not. This is the structural distinction that allows this new model of computation to escape the 1956 result.
4 Incomputability
Lemma 4.1 of [6] is a fundamental lemma for proving incomputability.
Lemma 4.1 (Fiske 2012).
Let $\varphi:\mathbb{N}\to\{0,1\}$ be an incomputable function. For each $k$ with $1\leq k\leq m$, let $B_{k}:\{0,1\}^{n}\to\{0,1\}^{n}$ be an invertible boolean function, and let $h:\mathbb{N}\to\{1,\ldots,m\}$ be a computable function. Define $g:\mathbb{N}\to\{0,1\}$ by
$\bigl(g(jn+1),\,\ldots,\,g((j+1)n)\bigr)\;=\;B_{h(j)}\bigl(\varphi(jn+1),\,\ldots,\,\varphi((j+1)n)\bigr).$
Then $g$ is incomputable.
The proof is by contradiction: if $g$ were computable, then since each $B_{k}$ is invertible and $h$ is computable, one could recover $\varphi$ by applying $B_{h(j)}^{-1}$ to each $n$-tuple, contradicting the incomputability of $\varphi$.
The dynamical level set construction generates an incomputable firing pattern sequence whenever the source of randomness measures quantum events. The Turing incomputability of the observable behavior is a mathematical consequence of the new level set concept, not an additional assumption.
5 Connection to the Principle of Self-Modifiability
This concept links the 2012 Turing Centenary paper [6] and subsequent papers on self-modifiable dynamical systems [7, 8]. In the AEM executing the UTM:
  • The state space is fixed: the set of UTM configurations.
  • The logical transition function $\eta$ is fixed: it is a specific UTM program.
  • The physical realization of the level sets of $\eta$ – the actual dynamical objects computing membership in $\eta_{k}^{-1}\{0\}$ and $\eta_{k}^{-1}\{1\}$ – are self-modifying during execution.
This construction illustrates the Principle of Self-Modifiability [7] applied to the level set structure of the system. The system modifies the geometric objects that define how it computes, while it is computing. The level sets as logical objects are invariant; the level sets as physical dynamical objects are self-modifiable. This distinction between the logical and physical identity of a mathematical structure is itself a new conceptual contribution.
6 Level sets and Phase spaces
The phase space framework inherited from Boltzmann [3], Poincaré [13], and Gibbs [11] treats every geometric object in the phase portrait – including level sets of the Lyapunov function [12], and the stable and unstable manifolds identified in Smale [15] – as a permanent feature of the landscape. The bifurcation theory of the Russian school [1, 2] allows the topology of these objects to change, but only at isolated points and always according to equations fixed in advance. None of this allows the physical realization of level sets to be a dynamical variable of the system itself.
The closest standard concepts are mathematically distinct:
  • Non-autonomous systems: the vector field depends on time, but according to a rule fixed before evolution begins.
  • Variable structure control: the controller switches between pre-specified rule sets, but the full menu of rules is fixed before execution.
  • Randomized Turing machines: random input does not change program structure, as established in the 1956 result.
Dynamic level sets do not appear to have been formalized, or studied in the standard mathematical literature prior to the paper, Turing Incomputable Computation.
7 Summary
The construction in [6], implemented as Procedure 2, implicitly introduces a new mathematical object: a system in which the logical structure of level sets is invariant, but their physical realization is a dynamical variable, evolving under the self-modifiability of the meta command in the AEM and incomputable quantum randomness [16, 9]. This mathematical object:
  1. 1.
    Escapes the 1956 de Leeuw, et. al. results by using the mechanism of a Meta-command level set realization;
  2. 2.
    Generates Turing incomputable behavior from a finite program, as proved by Theorem 4.2 via Lemma 4.1 of [6];
  3. 3.
    Achieves Shannon perfect secrecy [14] of machine execution, as proved by Theorem 4.3 of [6]; and
  4. 4.
    Anticipates the Principle of Self-Modifiability developed in [7], providing its earliest and perhaps most mathematically explicit instance implemented with an active element machine [5].

### v2

@@
-Dynamic Level Sets
+arXiv:2602.22530v2[cs.CC] 28 Feb 2026
 Abstract
 A mathematical concept is identified and analyzed that is implicit in the 2012 paper Turing Incomputable Computation [7], presented at the Alan Turing Centenary Conference (Turing 100, Manchester).
-The concept, called dynamic level sets, is distinct from mathematical concepts in the standard literature on dynamical systems, topology, and computability theory.
-A new mathematical object is explained and why it may have escaped prior characterizations, including the classical result of de Leeuw, Moore, Shannon, and Shapiro [4] that probabilistic Turing machines compute no more than deterministic ones.
+The concept, called dynamic level sets, is distinct from classical level sets, the Osher–Sethian level set method, and other standard mathematical concepts in dynamical systems, topology, and computability theory.
+The paper explains this new mathematical object and why it may have escaped prior characterizations, including the classical result of de Leeuw, Moore, Shannon, and Shapiro [4] that probabilistic Turing machines (with bias $p$ where $p$ is Turing computable) compute no more than deterministic ones.
+A key mechanism underlying the concept is the Principle of Self-Modifiability, whereby the physical realization of an invariant logical level set is reconfigured at each computational step by an incomputable process.
 1 Level Sets in Pure Mathematics
@@
 Trajectories move through them, but the level sets do not move.
-The entire geometric theory of dynamical systems (invariant manifolds, stable and unstable manifolds, basins of attraction, separatrices) rests on this assumption, as codified in the foundational survey of Smale [15] and in the Russian school’s treatment of bifurcations [1, 2].
+The entire geometric theory of dynamical systems (invariant manifolds, stable and unstable manifolds, basins of attraction, separatrices) rests on this assumption, as covered by the foundational survey of Smale [18] and in the Russian school’s treatment of bifurcations [1, 2].
 In the latter, bifurcations alter the topology of level sets at isolated parameter values, but always according to equations fixed in advance; between bifurcation points the level set structure is entirely permanent.
 Remark 1.1.
-This static property is not a technical limitation, but a foundational assumption.
+The permanence of level sets in classical dynamics is not a technical limitation but a foundational assumption.
 A classical dynamical system is a triple $(T,X,\Phi)$ where $\Phi:X\times T\to X$ is a fixed evolution rule.
-The level sets of any function associated with the system are therefore determined once and for all when $\Phi$ is specified.
+The level sets of any function associated with the system are static geometric objects: they do not move, deform, merge, or split.
+Trajectories move through them; the level sets themselves are permanent structures in phase space.
+This is a stricter condition than merely saying the level sets are determined by $\Phi$.
+A level set can be fully determined by a fixed rule and still change in time — as in the Osher–Sethian method described below, where the governing PDE is fixed in advance yet the zero level set $\Gamma(t)$ moves.
+The distinguishing feature of the classical notion of level sets is that stasis is a stricter requirement than determinacy: having a fixed rule is not sufficient to assure that level sets are permanent geometric structures, as the Osher–Sethian method illustrates.
+The Osher–Sethian Level Set Method (1988)
+A conceptually related but distinct notion was introduced by Osher and Sethian [5] in the context of interface propagation and computational fluid dynamics.
+In their method, one introduces a function $\varphi:\mathbb{R}^{n}\times[0,\infty)\to\mathbb{R}$ and tracks the zero level set
+$\Gamma(t)\;=\;\bigl\{\,x\in\mathbb{R}^{n}\;\bigm|\;\varphi(x,t)=0\,\bigr\}$
+as a moving interface. The evolution of $\varphi$ is governed by the fixed PDE
+$\frac{\partial\varphi}{\partial t}+F|\nabla\varphi|=0,$
+where $F$ is the speed that is a function of the curvature $\kappa$ and $|\nabla\varphi|$ is the Euclidean norm of the gradient of $\varphi$.
+Note that the defining rule for the zero level set is specified once and for all in advance.
+A crucial point is the following.
+The rule governing $\varphi$ is fixed: it is a predetermined PDE at $t=0$, set before the evolution of the geometric dynamical system starts.
+What changes in time is the level set $\Gamma(t)$ itself – surfaces may merge, split, or otherwise deform, but always as the deterministic output of the fixed governing equation.
+In this sense the Osher–Sethian method preserves the classical assumption that the evolution rule is permanently fixed; the novelty lies in representing a moving interface implicitly through the changing zero level set of a fixed-rule function, rather than tracking the interface explicitly.
+The Osher–Sethian level set differs from dynamic level sets defined in Section 4 below, where it is the logical level set (as a mathematical set) that is invariant, while its physical instantiation carries no fixed representation – the representation is not determined in advance, but reconfigured at each step by an incomputable process that uses self-modifiability.
+2 The Principle of Self-Modifiability
+A computer program can be viewed as a discrete, autonomous dynamical system [8].
+A set of differential equations specifies a different type of dynamical system that can sometimes model analog machines [15].
+When a dynamical system that performs a task is changed so that the task is no longer adequately performed, it is desirable to have a process or mechanism to heal the system.
+This means that part of the system should detect a change and part of the system should self-modify the damaged dynamical system back to the original system that adequately performed the task.
+The capability of a dynamical system to change itself is called self-modifiability.
+This principle was explicitly introduced in [9].
+A more concrete way to think about self-modifiability is to view a dynamical system as a system that is governed by a collection of rules.
+In the case of differential equations, each equation is a rule; in a computational machine, each machine instruction is a rule.
+A dynamical system is self-modifiable if it can change its own rules: this means the dynamical system can add new rules to itself; it can replace current rules with new rules; or it can delete rules.
+Generally, a self-modifiable dynamical system is non-autonomous because the rules governing its behavior can change as a function of time.
+In [6], the active element machine (AEM) is a computational machine that can add new rules when a “dynamical event” occurs during the machine’s execution.
+The AEM uses “meta commands” to change its own rules.
+More generally, self-modifiability need not always be triggered by damage to a dynamical system: in the next section, we discuss how the AEM uses self-modifiability as a routine mechanism, reconfiguring the physical realization of level sets at every computational step.
 3 Level Sets in the Active Element Machine
-The Active Element Machine (AEM) [5], when executing a Universal Turing Machine (UTM) program $\eta$, physically realizes the boolean transition functions
+The Active Element Machine (AEM) [6], when executing a Universal Turing Machine (UTM) program $\eta$, physically realizes the boolean functions (representing machine instructions)
 $\eta_{k}:\{0,1\}^{3}\times\{0,1\}^{2}\;\longrightarrow\;\{0,1\},\quad k\in\{0,1,2,3,4,5\},$
@@
 A dynamic level set decomposition of $f$ is a family of invertible boolean functions $\{B_{s}\}_{s\in S}$, indexed by a state space $S$, together with an incomputable sequence $\omega:\mathbb{N}\to\{0,1\}^{n}$ and a computable function $h:\mathbb{N}\to S$, such that at step $j$ the physical encoding of $f^{-1}\{1\}$ is determined by $B_{h(j)}(\omega(j))$ rather than by any fixed representation.
-The level set as a mathematical set is invariant; its physical instantiation is a dynamical variable.
+The level set as a mathematical set is invariant; its physical instantiation has no fixed representation.
 Remark 4.1.
-This captures what is constructed in Procedure 2  [6].
+Definition 4.1 captures what is constructed in Procedure 2  [7].
 The logical level sets $\eta_{k}^{-1}\{1\}$ are invariant properties of the UTM program $\eta$.
-The physical realizations are dynamical variables of the AEM, evolving by a rule that depends on the current UTM instruction (computable via $h$) and on quantum random input (incomputable by the two quantum randomness axioms; see [16, 9]).
+The physical realizations carry no fixed representation within the AEM; each representation is reconfigured at each computational step by a rule that depends on the current UTM instruction (computable via $h$) and on quantum random input.
 A dynamic level set is novel for the following reasons.
@@
 In a parametric family $F(x;\,\theta(t))$, the level sets $F^{-1}(c;\,\theta(t))$ sweep through a predetermined sequence as $\theta$ varies.
-Here, no predetermined sequence exists: the quantum randomness axioms guarantee that the sequence of level set realizations is itself Turing incomputable, so it cannot have been fixed in advance.
+Here, no predetermined sequence exists: the quantum randomness assures that the sequence of level set realizations is itself Turing incomputable, so it cannot have been fixed in advance.
 It is not bifurcation theory.
@@
 It is not a probabilistic Turing machine in the sense of de Leeuw et al.
-The classical 1956 result shows that a Turing machine augmented with a random oracle with probability $p$, where $p$ is Turing computable, computes no more than a deterministic one.
+The classical 1956 result shows that a probabilistic Turing machine with bias $p$, where $p$ is Turing computable, computes no more than a deterministic one.
 That result applies because the Turing machine program, and therefore its level set structure, does not change in response to random input.
-In the AEM, the Meta command reconfigures the physical realization of the level sets at each step based on the random input.
+In the AEM, the Meta command reconfigures the physical realization of the level sets at each step based on the quantum random input.
 The program’s logical content is invariant; its physical encoding is not.
@@
 The proof is by contradiction: if $g$ were computable, then since each $B_{k}$ is invertible and $h$ is computable, one could recover $\varphi$ by applying $B_{h(j)}^{-1}$ to each $n$-tuple, contradicting the incomputability of $\varphi$.
-The dynamical level set construction generates an incomputable firing pattern sequence whenever the source of randomness measures quantum events.
+The dynamic level set construction generates an incomputable firing pattern sequence whenever the source of randomness measures quantum events.
 The Turing incomputability of the observable behavior is a mathematical consequence of the new level set concept, not an additional assumption.
@@
 The system modifies the geometric objects that define how it computes, while it is computing.
-The level sets as logical objects are invariant; the level sets as physical dynamical objects are self-modifiable.
-This distinction between the logical and physical identity of a mathematical structure is itself a new conceptual contribution.
+The level sets as logical objects are invariant; the level sets as physical objects are self-modifiable — their representation can change via the meta command and quantum randomness.
+This distinction between the logical and physical identity of a mathematical structure is itself a new concept.
 7 Level sets and Phase spaces
@@
 The bifurcation theory of the Russian school [1, 2] allows the topology of these objects to change, but only at isolated points and always according to equations fixed in advance.
-None of this allows the physical realization of level sets to be a dynamical variable of the system itself.
-The closest standard concepts are mathematically distinct:
+None of this allows the physical realization of level sets to lack a fixed representation.
+The closest standard concepts are mathematically distinct.
   • Non-autonomous systems: the vector field depends on time, but according to a rule fixed before evolution begins.
   • Variable structure control: the controller switches between pre-specified rule sets, but the full menu of rules is fixed before execution.
-  • Randomized Turing machines: random input does not change program structure, as established in the 1956 result.
-Dynamic level sets do not appear to have been formalized, or studied in the standard mathematical literature prior to the paper, Turing Incomputable Computation.
+  • Probabilistic Turing machines: random input does not change program structure, as established in the 1956 result.
+Dynamic level sets, as defined herein, do not appear to have been formalized or studied in the mathematical literature prior to the paper, “Turing Incomputable Computation”.
 8 Summary
-The construction in [6], implemented as Procedure 2, implicitly introduces a new mathematical object: a system in which the logical structure of level sets is invariant, but their physical realization is a dynamical variable, evolving under the self-modifiability of the meta command in the AEM and incomputable quantum randomness [16, 9].
+The construction in [7], implemented as Procedure 2, implicitly introduces a new mathematical object: a system in which the logical structure of level sets is invariant, but their physical realization has no fixed representation – it is reconfigured at each step by the self-modifiability of the meta command in the AEM and quantum randomness [19, 11].
 This mathematical object:
   1. 1.
-    Escapes the 1956 de Leeuw, et. al. results by using the mechanism of a Meta-command level set realization;
+    Escapes the 1956 de Leeuw et al. results by using the mechanism of a Meta-command level set realization;
   2. 2.
@@
   4. 4.
-    Anticipates the Principle of Self-Modifiability developed in [7], providing its earliest and perhaps most mathematically explicit instance implemented with an active element machine [5].
+    Provides the earliest and perhaps most mathematically explicit instance of the Principle of Self-Modifiability [9], implemented with an active element machine [6].